OpenStudy (anonymous):

Solve: |2t + 5| = 3 A. {1, 4} B. {-1, -4} C. {-1, 4} D. {1, -4}

hartnn (hartnn):

when |a| = b then a=b or a =-b so what about |2t+5| = 3 ??

OpenStudy (anonymous):

i'm confused

hartnn (hartnn):

like we get 2 equations, one with positive sign and one with negative sign on right side. so for |2t+5|=3 we would have 2t+5 = 3 OR 2t+5 = -3 can you solve these 2 simple linear equations ?

OpenStudy (anonymous):

t=−1 t=−4

hartnn (hartnn):

correct! thats your answer :)

OpenStudy (anonymous):

Thanks :)

hartnn (hartnn):

welcome ^_^